Whispering Company sells a machine for $6,720 with a 12-month warranty agreement that requires the company to replace all defective parts and to provide the repair labor at no cost to the customers. With sales being made evenly throughout the year, the company sells 620 machines in 2017 (warranty expense is incurred half in 2017 and half in 2018). As a result of product testing, the company estimates that the warranty cost is $363 per machine ($158 parts and $205 labor). Use "Inventory" account to record the warranty expense. Assuming that actual warranty costs are incurred exactly as estimated, what journal entries would be made relative to the following facts?

Part 4

What amount, if any, is disclosed in the balance sheet as a liability for future warranty costs as of December 31, 2017?

Warranty liability$ ???? (blank box)

please show work
Answer #1

Warranty liability = $112530

Wages and salary payable (2017)= 620*205*50%= 63550

Warranty expense to be charged against inventory(2017)= 620*158*50%= 48980

Warranty liability as of December31, 2017 = 63550+48980 = 112530